A storage tank is a right circular cylinder 20 ft long and 8 ft in diameter with

its axis horizontal. If the tank is half full of olive oil weighing 57 lb/ft3, find the

work done in emptying it through a pipe that runs from the bottom of the

tank to an outlet that is 6 ft above the top of the tank.


1
Expert's answer
2022-05-11T10:59:12-0400

The amount of energy transferred is known as work done. Also one can say that force is required to move the object in the direction of force, i.e.,

"W = F \cdot d"

Work done may be positive, negative or zero, it depends upon the angle between the force and the displacement.

Here, the length of the right circular cylinder is 20 ft.

"W=\int^b_a F(x)dx"

a=-4 ft=-1.2192 m

b=6=1.8288 m

"V=\pi r^2 \Delta x"

r=8/2=4 ft=1.2192 m

"V=1.49\pi\Delta x"

57 lb/fr3=913 kg/m3

"m=Vd=1.49\pi\Delta x 913=1360.4 \pi \Delta x"

"F=mg=1360.4\pi\Delta x 9.8=13331.6 \pi \Delta x"


"W=\int^{1.8288}_{-1.2192}13331.6 \pi dx=13331.6 x|^{1.8288}_{-1.2192}=13331.6 (1.8288+1.2192)=40634.7 J"